# Preferential Latent Space Models for Networks with Textual Edges

### v1

@@ 3 Estimation
Given network data $\mathcal{Y}=\{{\bm{Y}}_{ij}\}_{i,j\in[n]}$, where ${\bm{Y}}_{ij}\in[0,1]^{m_{ij}\times K}$, we aim to estimate ${\bm{a}}$, ${\bm{W}}$, ${\bm{U}}$. Under model (1), the negative log-likelihood function, up to a constant, can be written as
$\ell\left(\bm{a},\bm{W},{\bm{U}}\right)=-\sum_{k=1}^{K}\sum_{i<j}^{n}\sum_{l=1}^{m_{ij}}\frac{1}{m_{ij}}\left\{y^{(k)}_{ijl}\Theta^{(k)}_{ij}+\log\left(1-\psi\left(\Theta^{(k)}_{ij}\right)\right)\right\},$ (2)
@@
$\min_{{\bm{a}}\in\mathbb{R}^{n},\bm{W}\in\mathbb{R}^{n\times K}_{+},{\bm{U}}\in\mathbb{R}^{n\times d}}\ell(\bm{a},\bm{W},{\bm{U}})\text{, \quad subject to}\|{\bm{W}}\|_{0}\leq s,$ (3)
where $s$ is a tuning parameter that controls the sparsity of ${\bm{W}}$. To enforce sparsity and the positivity constraint on ${\bm{W}}$ along the solution path, we employ a truncation operator Truncate(${\bm{W}},s$) defined as,
$[\operatorname{Truncate}({\bm{W}},s)]_{ik}=\begin{cases}W_{ik},&\text{ if }(i,k)\in\text{Supp}^{+}({\bm{W}},s),\\ 0,&\text{ otherwise, }\end{cases}$
for ${\bm{W}}\in\mathbb{R}^{n\times K}$ and $s\leq nK$. The set $\text{Supp}^{+}({\bm{W}},s)$ denotes the $s$ entries in ${\bm{W}}$ with the largest values. To solve (3), we develop a projected gradient descent algorithm that is easy to implement and computationally efficient. Our estimation procedure is summarized in Algorithm 1.
@@
  Input: network data $\mathcal{Y}$, initial values ${\bm{a}}^{(0)},\bm{W}^{(0)},{\bm{U}}^{(0)}$, step sizes $\eta_{{\bm{a}}},\eta_{{\bm{W}}},\eta_{{\bm{U}}}$.
  repeat for $t=0,1,\ldots,$     ${\bm{a}}^{(t+1)}={\bm{a}}^{(t)}-\eta_{{\bm{a}}}\nabla_{{\bm{a}}}\ell({\bm{a}},{{\bm{W}}}^{(t)},{\bm{U}}^{(t)})|_{{\bm{a}}={\bm{a}}^{(t)}}$;     ${{\bm{W}}}^{(t+1)}={\rm Truncate}\left[{\bm{W}}^{(t)}-\eta_{{\bm{W}}}\nabla_{{\bm{W}}}\ell({\bm{a}}^{(t)},{\bm{W}},{\bm{U}}^{(t)})|_{{\bm{W}}={\bm{W}}^{(t)}},s\right]$;     ${{\bm{U}}}^{(t+1)}={\bm{U}}^{(t)}-\eta_{{\bm{U}}}\nabla_{{\bm{U}}}\ell({\bm{a}}^{(t)},{{\bm{W}}}^{(t)},{\bm{U}})|_{{\bm{U}}={\bm{U}}^{(t)}}$; normalize rows of ${{\bm{U}}}^{(t+1)}$;
@@
The parameters $\eta_{{\bm{a}}},\eta_{{\bm{W}}},\eta_{{\bm{U}}}$ control the step sizes in the gradient descent algorithm. Theorem 1 provides theoretical conditions on $\eta_{{\bm{a}}},\eta_{{\bm{W}}},\eta_{{\bm{U}}}$ to ensure the algorithm achieves a linear convergence rate. Based on these conditions, we discuss practical choices for these parameters. In practice, backtracking line search can be implemented for $\eta_{{\bm{a}}},\eta_{{\bm{W}}},\eta_{{\bm{U}}}$ at each step of the iteration to achieve fast convergence.
@@ 4 Theoretical Results
$\mathcal{F}_{n,d,k}(M_{1})=\bigg{\{}({\bm{a}},{\bm{W}},{\bm{U}})\bigg{|}$ $\,\,\|{\bm{a}}\|_{\infty}\leq\frac{M_{1}}{4},\,\,\max_{i}\sum_{k}(W_{ik})^{2}\leq\frac{M_{1}}{2},\,\,\|{\bm{W}}\|_{0}<nk,$ (4)
$\|\bm{u}_{i}\|_{2}=1,\,\,\max_{i\neq j,k}\Theta_{ij}^{(k)}\leq-(1-C)M_{1}\bigg{\}},$
where $M_{1}\geq 0$ is a scalar that may depend on $n$ and $10<C<$ 1 is a constant. By the definition of $\Theta_{ij}^{(k)}$ in (1) and combining $\|{\bm{a}}\|_{\infty}\leq{M_{1}}/{4}$, $\max_{i}\sum_{k}(W_{ik})^{2}\leq{M_{1}}/{2}$ and $\|\bm{u}_{i}\|_{2}=1$ in (4), it is straightforward to show that $\max_{i,j,k}|\Theta_{ij}^{(k)}|\leq M_{1}$. Hence, for any $({\bm{a}},{\bm{W}},{\bm{U}})\in\mathcal{F}_{n,d,k}(M_{1})$, $\Theta_{ij}^{(k)}$’s are uniformly bounded in $[-M_{1},-(1-C)M_{1}]$ for any $k$ and $i\neq j$. That is, edge probabilities $\psi({\Theta_{ij}^{(k)}})$’s are bounded between $\frac{1}{1+e^{M_{1}}}$ and $\frac{1}{1+e^{(1-C)M_{1}}}$. It is seen that $M_{1}$ controls the overall sparsity of the network. If, for example, $M_{1}$ is in the order of $\log(n)-\log\log(n)$, then the average edge probability is in the order of $\log n/n$.
Let $({\bm{a}}^{\ast},{\bm{W}}^{\ast},{\bm{U}}^{\ast})$ be the true parameter, and write the nonzero singular values of ${\bm{U}}^{\ast}$ as $\sigma_{1}^{\ast}\geq\cdots\geq\sigma_{d}^{\ast}>0$ and $s^{\ast}=\|{\bm{W}}^{\ast}\|_{0}$. Write $w_{\max}=\max_{k}{{w^{(k)}_{\max}}}$, where ${{w^{(k)}_{\max}}}=\max_{i}W_{ik}^{\ast}$ is the maximum entry in column ${\bm{W}}^{\ast}_{.k}$, and $w_{\min}=\min_{k}{w^{(k)}_{\min}}$, where ${w^{(k)}_{\min}}=\min_{i:W_{ik}^{\ast}\neq 0}W_{ik}^{\ast}$ is the minimum nonzero entry in column ${\bm{W}}^{\ast}_{.k}$. We assume $w_{\max}\asymp{{w^{(k)}_{\max}}}$ and $w_{\min}\asymp{w^{(k)}_{\min}}$ for any $k$. This assumption is made to simplify notations in our analysis, and our results hold under more general conditions on $W_{ik}^{\ast}$’s but with more involved notations. We denote $\bar{m}=\left(\max_{i}\frac{1}{n}\sum_{j}\frac{1}{m_{ij}}\right)^{-1}$, where $\bar{m}$ characterizes the average number of edges. To simplify notation, we assume $\min_{ij}m_{ij}=O(1)$, that is, the minimum number of edges between two nodes is a constant.
To investigate the computational and statistical properties of iterates from Algorithm 1, we first introduce an error metric. As ${\bm{U}}$ is identifiable up to an orthogonal transformation, for any ${\bm{U}}_{1},{\bm{U}}_{2}\in\mathbb{R}^{n\times d}$, we define a distance measure
$\text{dist}({\bm{U}}_{1},{\bm{U}}_{2})=\min_{{\bm{R}}:{\bm{R}}{\bm{R}}^{\top}={\bm{I}}}\|{\bm{U}}_{1}-{\bm{U}}_{2}{\bm{R}}\|_{F}.$
@@
$e_{t}=2Kn\|{\bm{a}}^{(t)}-{\bm{a}}^{\ast}\|_{2}^{2}+{\sigma_{1}^{\ast}}^{2}w_{\max}^{2}\|{\bm{W}}^{(t)}-{\bm{W}}^{\ast}\|_{F}^{2}+K{\sigma_{1}^{\ast}}^{2}w_{\max}^{4}\text{dist}^{2}({\bm{U}}^{(t)},{\bm{U}}^{\ast}).$ (5)
We first derive an error bound for $e_{t}$ in Theorem 1, and then derive error bounds for ${\bm{a}}^{(t)}$, ${\bm{W}}^{(t)}$ and ${\bm{U}}^{(t)}$, respectively, in Corollary 1. We assume the following regularity conditions.
@@
Let $\kappa_{0}=\frac{{\sigma_{1}^{\ast}}w_{\max}^{2}}{{\sigma_{d}^{\ast}}w_{\min}^{2}}$. Assume initial values ${\bm{a}}^{(0)}$, ${\bm{W}}^{(0)}$ and ${\bm{U}}^{(0)}$ satisfy
@@
This assumption requires the initial values to be reasonably close to the true parameters. Such assumptions are commonly employed in nonconvex optimizations (Lyu et al., 2023; Zhang et al., 2023). In particular, if $\kappa_{0}=O(1)$ and $d=O(1)$, then Assumption 1 can be simplified to $\|{\bm{a}}^{(0)}-{\bm{a}}^{\ast}\|_{2}^{2}=O(nw_{\max}^{4}e^{-2M_{1}})$, $\|{\bm{W}}^{(0)}-{\bm{W}}^{\ast}\|_{F}^{2}=O(Knw_{\max}^{2}e^{-2M_{1}})$ and $\text{dist}^{2}({\bm{U}}^{(0)},{\bm{U}}^{\ast})=O(ne^{-2M_{1}})$. These assumptions on ${\bm{a}}^{(0)}$, ${\bm{W}}^{(0)}$ and ${\bm{U}}^{(0)}$ are relatively mild.
@@
$K{\sigma_{d}^{\ast}}^{2}\geq C_{2}\frac{w_{\max}^{2}}{w_{\min}^{4}}\max\left\{\frac{n}{\bar{m}},\log(n)\right\}e^{CM_{1}}.$
This is an assumption on the minimal signal strength $\sigma_{d}^{\ast}$, which is the minimum nonzero singular value of ${\bm{U}}^{\ast}$. It is seen that the signal strength condition weakens as the number of layers $K$ or average number of edges $\bar{m}$ increases. Also, the signal strength condition becomes stronger as $M_{1}$ increases, corresponding to sparser networks.
Next, we are ready to state our main theorem.
@@
Suppose $({\bm{a}}^{(0)},{\bm{W}}^{(0)},{\bm{U}}^{(0)})$ satisfies Assumption 1 and $({\bm{a}}^{\ast},{\bm{W}}^{\ast},{\bm{U}}^{\ast})$ is in (4) and satisfies Assumption 2. Letting $\eta_{\bm{a}}=\frac{\eta}{4Kn}$, $\eta_{{\bm{W}}}=\frac{\eta}{4{\sigma_{1}^{\ast}}^{2}w_{\max}^{2}}$, $\eta_{{\bm{U}}}=\frac{\eta}{2K{\sigma_{1}^{\ast}}^{2}w_{\max}^{4}}$ and $s=\gamma s^{\ast}$ for $\gamma>1$, the $t$-th step iteration of Algorithm 1 satisfies, with probability as least $1-Kn^{-1}$,
$e_{t}\precsim\rho^{t}e_{0}+\kappa_{0}^{4}e^{(1+C)M_{1}}\left\{d\times\max\left\{\frac{n}{\bar{m}},\log(n)\right\}+\frac{s^{\ast}\log(n)}{\bar{m}}\right\},$
where $0<\rho<\frac{1}{2}$ and $\eta=\frac{1}{4}\kappa_{0}^{2}(16-\rho)e^{M_{1}}$.
@@
This theorem describes the estimation error at each iteration and provides theoretical guidance on the step sizes $\eta_{{\bm{a}}}$, $\eta_{{\bm{W}}}$ and $\eta_{\bm{U}}$ used in Algorithm 1. The error bound in Theorem 1 consists of two terms. The first term, $\rho^{t}e_{0}$, is the computational error, which decays geometrically with the iteration number $t$ since the contraction parameter $\rho$ satisfies $0<\rho<\frac{1}{2}$. The second term, $\kappa_{0}^{4}e^{(1+C)M_{1}}\left\{d\times\max\left\{\frac{n}{\bar{m}},\log(n)\right\}+\frac{s^{\ast}\log(n)}{\bar{m}}\right\}$, represents the statistical error, which is related to the stochasticity in the data and does not vary with the iteration number $t$. The error bound in Theorem 1 reveals an interesting interplay between the computational efficiency and statistical rate of convergence. Specifically, when the number of iterations is sufficiently large, the computational error is to be dominated by the statistical error and the resulting estimator falls within the statistical precision of the true parameters. In the statistical error, the term $\frac{s^{\ast}\log(n)}{\bar{m}}$ is related to estimating the sparse matrix ${\bm{W}}^{\ast}$ and the term $d\times\max\left\{\frac{n}{\bar{m}},\log(n)\right\}$ is related to estimating the low-rank matrix ${\bm{U}}^{\ast}$. The statistical error decreases with the average number of edges $\bar{m}$ and increases with the sparsity parameter $M_{1}$. We note that, by the definition in (5), the error metric $e_{t}$ increases with $n$ and $K$. The estimation errors of ${\bm{a}}^{\ast}$, ${\bm{W}}^{\ast}$ and ${\bm{U}}^{\ast}$ are provided in Corollary 1.
@@
Our theoretical analysis is nontrivial. One key component of our analysis is a tight concentration bound on the spectrum of a random matrix, where the $(i,j)$th entry is an average of $m_{ij}$ binary random variables. This bound, detailed in Lemma S5 in the supplement, is sharper than the matrix Bernstein inequality (Tropp, 2012) and can accommodate varying $m_{ij}$ across $i,j\in[n]$. The proof of this concentration result involves intricate technical details, and it uses large deviation estimates and geometric functional analysis techniques. It extends the result in Lei and Rinaldo (2015), which was derived for the case of $m_{ij}=1$. Using Lemma S5 we are able to improve the statistical error for low rank matrix ${\bm{U}}^{\ast}$ from $\kappa_{0}^{4}e^{(1+C)M_{1}}dn\log(n)$, which can be derived using the matrix Bernstein inequality (Tropp, 2012) under $m_{ij}=1$, to $\kappa_{0}^{4}e^{(1+C)M_{1}}d\times\max\left\{\frac{n}{\bar{m}},\log(n)\right\}$, which in turn relaxes the minimal signal strength condition in Assumption 2 and highlights the benefit of having a greater average number of edges $\bar{m}$. Additionally, introducing sparse preferential effects in the latent space model further complicates our theoretical analysis due to the interplay between ${\bm{W}}$ and ${\bm{U}}$. This requires carefully bounding the error of ${\bm{W}}^{(t)}$ and ${\bm{U}}^{(t)}$ in each step of the iteration to achieve contraction while ensuring the identifiability conditions are met.
@@
$t\geq\log\left(\frac{d\times\max\{\frac{n}{\bar{m}},\log(n)\}+\frac{s^{\ast}\log(n)}{\bar{m}}}{C_{1}K{\sigma_{1}^{\ast}}^{4}w_{\max}^{4}}\kappa_{0}^{8}e^{(3+C)M_{1}}\right)/\log(\rho)$, it holds that
$\|{\bm{a}}^{(t)}-{\bm{a}}^{\ast}\|_{2}^{2}\precsim\frac{\kappa_{0}^{4}e^{(1+C)M_{1}}}{K}\left\{d\times\max\left\{\frac{1}{\bar{m}},\frac{\log(n)}{n}\right\}+\frac{s^{\ast}\log(n)}{n\bar{m}}\right\},$
$\|{\bm{W}}^{(t)}-{\bm{W}}^{\ast}\|_{F}^{2}\precsim\frac{\kappa_{0}^{4}e^{(1+C)M_{1}}}{w_{\max}^{2}}\left\{d\times\max\left\{\frac{1}{\bar{m}},\frac{\log(n)}{n}\right\}+\frac{s^{\ast}\log(n)}{n\bar{m}}\right\},$
$\text{dist}^{2}({\bm{U}}^{(t)},{\bm{U}}^{\ast})\precsim\frac{\kappa_{0}^{4}e^{(1+C)M_{1}}}{Kw_{\max}^{4}}\left\{d\times\max\left\{\frac{1}{\bar{m}},\frac{\log(n)}{n}\right\}+\frac{s^{\ast}\log(n)}{n\bar{m}}\right\},$
@@
This corollary establishes upper bounds on the estimation errors for ${\bm{a}}^{(t)}$, ${\bm{W}}^{(t)}$, and ${\bm{U}}^{(t)}$, respectively. Notably, the error bounds for ${\bm{a}}^{(t)}$ and ${\bm{U}}^{(t)}$ decrease with $K$, indicating that their estimation improves as the number of layers $K$ increases. All three estimation errors decrease with the average number of edges $\bar{m}$, suggesting that observing more edges between nodes leads to better estimation. Finally, the estimation error for ${\bm{U}}^{\ast}$ aligns with that in standard latent space models (Ma et al., 2020) when $K=1$, $\bar{m}=1$ and $W_{ij}=1$ for all $i,j$.

### v2

@@ 3 Estimation
Given network data $Y=\{{Y}_{ij}\}_{i,j\in[n]}$, where ${Y}_{ij}\in[0,1]^{m_{ij}\times K}$, we aim to estimate ${a}$, ${W}$, and ${U}$. Under model (1), the negative log-likelihood function, up to a constant, can be written as
$\ell\left(a,{W},{U}\right)=-\sum_{k=1}^{K}\sum_{i<j}^{n}\sum_{l=1}^{m_{ij}}\frac{1}{m_{ij}}\left\{y^{(k)}_{ijl}\Theta^{(k)}_{ij}+\log\left(1-\psi(\Theta^{(k)}_{ij})\right)\right\},$ (2)
@@
$\min_{{a}\in\mathbb{R}^{n},{W}\in\mathbb{R}^{n\times K}_{+},{U}\in\mathbb{R}^{n\times d}}\ell(a,{W},{U})\text{, \quad subject to }\|{W}\|_{0}\leq s,$ (3)
where $s$ is a tuning parameter that controls the sparsity of ${W}$. To enforce sparsity and the positivity constraint on ${W}$ along the solution path, we employ a truncation operator Truncate(${W},s$) for $s\leq nK$ such that $[\operatorname{Truncate}({W},s)]_{ik}=W_{ik}$ if $(i,k)\in\text{Supp}^{+}({W},s)$ and $[\operatorname{Truncate}({W},s)]_{ik}=0$ otherwise, where $\text{Supp}^{+}({W},s)$ denotes the $s$ entries in ${W}$ with the largest values. To solve (3), we develop a projected gradient descent algorithm that is easy to implement and computationally efficient. Our estimation procedure is summarized in Algorithm 1.
@@
  Input: network data $Y$, initial values ${a}^{(0)},\bm{W}^{(0)},{U}^{(0)}$, step sizes $\eta_{{a}},\eta_{{W}},\eta_{{U}}$.
  repeat for $t=0,1,\ldots,$     ${a}^{(t+1)}={a}^{(t)}-\eta_{{a}}\nabla_{{a}}\ell({a},{{W}}^{(t)},{U}^{(t)})|_{{a}={a}^{(t)}}$;     ${{W}}^{(t+1)}={\rm Truncate}\left({W}^{(t)}-\eta_{{W}}\nabla_{{W}}\ell({a}^{(t)},{W},{U}^{(t)})|_{{W}={W}^{(t)}},s\right)$;     ${{U}}^{(t+1)}={U}^{(t)}-\eta_{{U}}\nabla_{{U}}\ell({a}^{(t)},{{W}}^{(t)},{U})|_{{U}={U}^{(t)}}$; normalize rows of ${{U}}^{(t+1)}$;
@@
  Output $a$, $W$, and $U$
The parameters $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ control the step sizes in the gradient descent algorithm. Theorem 1 provides theoretical conditions on $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ to ensure the algorithm achieves a linear convergence rate. In practice, backtracking line search can be implemented for $\eta_{{a}},\eta_{{W}},\eta_{{U}}$ at each step of the iteration to achieve fast convergence.
@@ 4 Theoretical Results
$\mathcal{F}_{n,d,K}(M_{1})=\bigg{\{}({a},{W},{U})\mid$ $\,\,\|{a}\|_{\infty}\leq{M_{1}}/{4},\,\,\max_{i}\sum_{k=1}^{K}(W_{ik})^{2}\leq{M_{1}}/{2},\,\,\|{W}\|_{0}<nK,$ (4)
$\|u_{i}\|_{2}=1,\,\,\max_{i\neq j,k}\Theta_{ij}^{(k)}\leq-(1-C)M_{1}\bigg{\}},$
where $M_{1}\geq 0$ is a scalar that may depend on $n$ and $10<C<$ 1 is a constant. By the definition of $\Theta_{ij}^{(k)}$ in (1) and combining $\|{a}\|_{\infty}\leq{M_{1}}/{4}$, $\max_{i}\sum_{k=1}^{K}(W_{ik})^{2}\leq{M_{1}}/{2}$ and $\|u_{i}\|_{2}=1$ in (4), it is straightforward to show that $\max_{i,j,k}|\Theta_{ij}^{(k)}|\leq M_{1}$. Hence, for any $({a},{W},{U})\in\mathcal{F}_{n,d,K}(M_{1})$, $\Theta_{ij}^{(k)}$’s are uniformly bounded in $[-M_{1},-(1-C)M_{1}]$ for any $k$ and $i\neq j$. That is, edge probabilities $\psi({\Theta_{ij}^{(k)}})$’s are bounded between ${1}/({1+e^{M_{1}}})$ and ${1}/({1+e^{(1-C)M_{1}}})$. It is seen that $M_{1}$ controls the overall sparsity of the network. If, for example, $M_{1}$ is in the order of $\log(n)-\log\log(n)$, then the average edge probability is in the order of $\log(n)/n$.
Let $({a}^{\ast},{W}^{\ast},{U}^{\ast})$ be the true parameter, $\sigma_{1}^{\ast}\geq\cdots\geq\sigma_{d}^{\ast}>0$ be the nonzero singular values of ${U}^{\ast}$ and $s^{\ast}=\|{W}^{\ast}\|_{0}$. Write $w_{\max}=\max_{k}{{w^{(k)}_{\max}}}$, where ${{w^{(k)}_{\max}}}=\max_{i}W_{ik}^{\ast}$ is the maximum entry in column ${W}^{\ast}_{.k}$, and $w_{\min}=\min_{k}{w^{(k)}_{\min}}$, where ${w^{(k)}_{\min}}=\min_{i:W_{ik}^{\ast}\neq 0}W_{ik}^{\ast}$ is the minimum nonzero entry in column ${W}^{\ast}_{.k}$. We assume $w_{\max}\asymp{{w^{(k)}_{\max}}}$ and $w_{\min}\asymp{w^{(k)}_{\min}}$ for any $k$. This assumption is made to simplify notations in our analysis, and our results hold under more general conditions on $W_{ik}^{\ast}$’s but with more involved notations. We denote $\bar{m}=(\max_{i}{1}/{n}\sum_{j}{1}/{m_{ij}})^{-1}$, where $\bar{m}$ characterizes the average number of edges. To further simplify notation, we assume $\min_{ij}m_{ij}=O(1)$, that is, the minimum number of edges between two nodes is a constant.
We first introduce an error metric for the iterates from Algorithm 1. As ${U}$ is identifiable up to an orthogonal transformation, for any ${U}_{1},{U}_{2}\in\mathbb{R}^{n\times d}$, we define a distance measure
$\text{dist}({U}_{1},{U}_{2})=\min_{{R}:{R}{R}^{\top}={I}}\|{U}_{1}-{U}_{2}{R}\|_{F}.$
@@
$e_{t}=2Kn\|{a}^{(t)}-{a}^{\ast}\|_{2}^{2}+{\sigma_{1}^{\ast}}^{2}w_{\max}^{2}\|{W}^{(t)}-{W}^{\ast}\|_{F}^{2}+K{\sigma_{1}^{\ast}}^{2}w_{\max}^{4}\text{dist}^{2}({U}^{(t)},{U}^{\ast}).$ (5)
We first derive an error bound for $e_{t}$ in Theorem 1, and then derive error bounds for ${a}^{(t)}$, ${W}^{(t)}$ and ${U}^{(t)}$, respectively, in Corollary 1. We assume the following regularity conditions.
@@
Let $\kappa_{0}=({{\sigma_{1}^{\ast}}w_{\max}^{2}})/({{\sigma_{d}^{\ast}}w_{\min}^{2}})$. Assume initial values ${a}^{(0)}$, ${W}^{(0)}$ and ${U}^{(0)}$ satisfy
@@
This assumption requires the initial values to be reasonably close to the true parameters. Such assumptions are commonly employed in nonconvex optimizations (Lyu et al., 2023; Zhang et al., 2023). In particular, if $\kappa_{0}=O(1)$ and $d=O(1)$, then Assumption 1 can be simplified to $\|{a}^{(0)}-{a}^{\ast}\|_{2}^{2}=O(nw_{\max}^{4}e^{-2M_{1}})$, $\|{W}^{(0)}-{W}^{\ast}\|_{F}^{2}=O(Knw_{\max}^{2}e^{-2M_{1}})$ and $\text{dist}^{2}({U}^{(0)},{U}^{\ast})=O(ne^{-2M_{1}})$. These assumptions on ${a}^{(0)}$, ${W}^{(0)}$ and ${U}^{(0)}$ are relatively mild.
@@
$K{\sigma_{d}^{\ast}}^{2}\geq C_{2}({w_{\max}^{2}}/{w_{\min}^{4}})\max\left\{{n}/{\bar{m}},\log(n)\right\}e^{CM_{1}}.$
This is an assumption on the minimal signal strength $\sigma_{d}^{\ast}$, which is the minimum nonzero singular value of ${U}^{\ast}$. It is seen that the signal strength condition weakens as the number of layers $K$ or average number of edges $\bar{m}$ increases. Also, the signal strength condition becomes stronger as $M_{1}$ increases, corresponding to sparser networks.
@@
Suppose $({a}^{(0)},{W}^{(0)},{U}^{(0)})$ satisfies Assumption 1 and $({a}^{\ast},{W}^{\ast},{U}^{\ast})$ is in (4) and satisfies Assumption 2. Letting $\eta_{a}={\eta}/({4Kn})$, $\eta_{{W}}={\eta}/({4{\sigma_{1}^{\ast}}^{2}w_{\max}^{2}})$, $\eta_{{U}}={\eta}/({2K{\sigma_{1}^{\ast}}^{2}w_{\max}^{4}})$ and $s=\gamma s^{\ast}$ for $\gamma>1$, the $t$-th step iteration of Algorithm 1 satisfies, with probability as least $1-Kn^{-1}$,
$e_{t}\precsim\rho^{t}e_{0}+\kappa_{0}^{4}e^{(1+C)M_{1}}\left[d\max\left\{{n}/{\bar{m}},\log(n)\right\}+{s^{\ast}\log(n)}/{\bar{m}}\right],$
where $0<\rho<{1}/{2}$ and $\eta=\kappa_{0}^{2}(16-\rho)e^{M_{1}}/{4}$.
@@
This theorem describes the estimation error at each iteration and provides theoretical guidance on step sizes $\eta_{{a}}$, $\eta_{{W}}$ and $\eta_{U}$ in Algorithm 1. The error bound consists of two terms. The first term $\rho^{t}e_{0}$ is the computational error, which decays geometrically with the iteration number $t$ since the contraction parameter $\rho$ satisfies $0<\rho<{1}/{2}$. The second term $\kappa_{0}^{4}e^{(1+C)M_{1}}\left[d\max\left\{{n}/{\bar{m}},\log(n)\right\}+{s^{\ast}\log(n)}/{\bar{m}}\right]$ represents the statistical error, which is related to noise in the data and does not vary with $t$. These two terms reveal an interesting interplay between the computational efficiency and statistical rate of convergence. Specifically, when the number of iterations is sufficiently large, the computational error is to be dominated by the statistical error and the resulting estimator falls within the statistical precision of the true parameters. In the statistical error, the term ${s^{\ast}\log(n)}/{\bar{m}}$ is related to estimating the sparse matrix ${W}^{\ast}$ and the term $d\max\left\{{n}/{\bar{m}},\log(n)\right\}$ is related to estimating the low-rank matrix ${U}^{\ast}$. The statistical error decreases with the average number of edges $\bar{m}$ and increases with the sparsity parameter $M_{1}$.
@@
Our theoretical analysis is nontrivial as the node-layer preferential effects in $W_{ik}W_{jk}u_{i}u_{j}$ lead to an involved interplay between ${W}$ and ${U}$. This requires carefully bounding the error of ${W}^{(t)}$ and ${U}^{(t)}$ (up to rotation) in each step of the iteration to achieve contraction while ensuring the identifiability conditions are met. To tackle varying edge number $m_{ij}$’s, we derive a tight bound on the spectrum of random matrices with bounded moments following the techniques in Bandeira and van Handel (2016); see Lemma S5. This bound is sharper than the matrix Bernstein inequality (Tropp, 2012). Using Lemma S5, we are able to improve the statistical error for low rank matrix ${U}^{\ast}$ from $\kappa_{0}^{4}e^{(1+C)M_{1}}dn\log(n)$, which can be derived using the matrix Bernstein inequality under $m_{ij}=1$, to $\kappa_{0}^{4}e^{(1+C)M_{1}}d\max\left\{{n}/{\bar{m}},\log(n)\right\}$, which in turn relaxes the minimal signal strength condition in Assumption 2.
@@
$t\geq\log\left[\{{d\max\{{n}/{\bar{m}},\log(n)\}+{s^{\ast}\log(n)}/{\bar{m}}}\}\kappa_{0}^{8}e^{(3+C)M_{1}}/({C_{1}K{\sigma_{1}^{\ast}}^{4}w_{\max}^{4}})\right]/\log(\rho),$ it holds that
$\|{a}^{(t)}-{a}^{\ast}\|_{2}^{2}\precsim\frac{\kappa_{0}^{4}e^{(1+C)M_{1}}}{K}\left[d\max\left\{\frac{1}{\bar{m}},\frac{\log(n)}{n}\right\}+\frac{s^{\ast}\log(n)}{n\bar{m}}\right],$
$\|{W}^{(t)}-{W}^{\ast}\|_{F}^{2}\precsim\frac{\kappa_{0}^{4}e^{(1+C)M_{1}}}{w_{\max}^{2}}\left[d\max\left\{\frac{1}{\bar{m}},\frac{\log(n)}{n}\right\}+\frac{s^{\ast}\log(n)}{n\bar{m}}\right],$
$\text{dist}^{2}({U}^{(t)},{U}^{\ast})\precsim\frac{\kappa_{0}^{4}e^{(1+C)M_{1}}}{Kw_{\max}^{4}}\left[d\max\left\{\frac{1}{\bar{m}},\frac{\log(n)}{n}\right\}+\frac{s^{\ast}\log(n)}{n\bar{m}}\right],$
@@
In Corollary 1, the error bounds for ${a}^{(t)}$ and ${U}^{(t)}$ decrease with $K$, indicating that their estimation improves as the number of layers $K$ increases. All three estimation errors decrease with the average number of edges $\bar{m}$, suggesting that observing more edges between nodes leads to better estimation. Finally, the estimation error for ${U}^{\ast}$ matches with that in standard latent space models (Ma et al., 2020) when $K=1$, $\bar{m}=1$ and $W_{ij}=1$ for all $i,j$.